OpenStudy (anonymous):

identify the equation as either an identity or not plzz help meeeeeee sin x/1-cos x+sin x/1+cos x=2csc x

OpenStudy (anonymous):

cos2x=1-2sin^2x=2cos^x-1 use them for 1-cosx and 1+cosx

OpenStudy (anonymous):

can u help me plz i didnt get get at all

OpenStudy (anonymous):

cox=1-2sin^2(x/2) 1-cosx =2sin^x/2 1+cosx=2cos^2x/2

OpenStudy (anonymous):

or u could solve normally sinx/1-cosx + sinx/1+cosx =sinx(1-cosx+1+cosx)/1-cos^2x =2sinx/sin^2x =2/sinx =2cosecx this method is quick i guesss

OpenStudy (anonymous):

can u write it on the board so i can see it plz if u dont mind i t

OpenStudy (anonymous):



OpenStudy (anonymous):

so is it identity or not

OpenStudy (anonymous):

so is it identity or not

OpenStudy (anonymous):

it is identity

OpenStudy (anonymous):

oh ok thanks bro

OpenStudy (anonymous):

u r welcome